Datacenter Dallas Texas on 28 Dec 2025 19:21:46 (UTC-05:00)
Resolved on 28 Dec 2025 20:57:11 (UTC-05:00)
All services are restored.
In progress on 28 Dec 2025 19:55:13 (UTC-05:00)
The facility has reached approximately 80* F, we will begin slowly powering services back up.
In progress on 28 Dec 2025 19:28:12 (UTC-05:00)
The facility is cooling off, but it is still too warm to begin powering equipment back up. We estimate approximately 15 more minutes until the facility reaches acceptable temperatures to begin powering equipment back up.
Identified on 28 Dec 2025 19:23:17 (UTC-05:00)
Facilities personnel have arrived at the datacenter, and CRAC units and additional rooftop units are being re-energized. It will take approximately 30 minutes to cool the facility to acceptable levels for servers to power up without damage.
Investigating on 28 Dec 2025 19:22:37 (UTC-05:00)
We are seeing high temperatures across many devices that are still online. We have notified our provider
Investigating on 28 Dec 2025 19:22:11 (UTC-05:00)
We are seeing high temperatures on core network devices; we are continuing to investigate.
Investigating on 28 Dec 2025 19:21:46 (UTC-05:00)
At approximately 15:06 Pacific time, our NOC began to see servers going offline.